OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (3, 2) and (5, 12)

OpenStudy (mathstudent55):

1. Choose one point as first point and other point as second point. 2. Subtract first y value from second y value 3. Subtract first x value from second x value 4. Divide the difference of y values by the difference of the x values That is what is meant by the formula for the slope: slope = m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

so would the slope be -1/5

OpenStudy (anonymous):

\[\frac{ (12-2) }{ (5-3) }\]

OpenStudy (anonymous):

10/3

OpenStudy (anonymous):

5-3=2 10/2 = 5